OpenStudy (anonymous):

What is the maximum number of possible solutions for the system shown below? x^2-4y^2=64 x^2+y^2=36 A. 2 B. 3 C. 1 D. 4

OpenStudy (anonymous):

Ellipse + hyperbola max 4

OpenStudy (dumbcow):

well \[y = \pm \sqrt{28/3}\] \[x = \pm \sqrt{304/3}\] how many is that?

OpenStudy (anonymous):

four

OpenStudy (dumbcow):

yep :)

OpenStudy (anonymous):

but they have to be pairs to be called solution so 2

OpenStudy (dumbcow):

(x,y) (-x,y) (x,-y) (-x,-y) All solutions because the variable is being squared in the equations

OpenStudy (anonymous):

i don't remenber it now but last year when we coped with these systems only quite a few had 4 solutions....

OpenStudy (dumbcow):

or look at estudiar answer if you draw an ellipse and a hyperbola, that can cross at 4 points